Through this chapter,you will be using and comparing fractions and decimals.Cicle all of the fractions below that are equal to 1/2.
6/12 15/30 4/7 5/10 1/3 2/3 9/12 7/14
1
Expert's answer
2011-10-07T11:26:44-0400
The rule we may use is
(ac)/(bc) = a/b (for “c” that isn’t zero)
So, let’s start.
6/12 = (6x1)/(6x2) = 1/2
15/30 = (15x1)/(15x2) = 1/2
4/7 isn’t equal to 1/2
5/10 = (5x1)/(5x2) = 1/2
1/3 isn’t equal to 1/2
2/3 isn’t equal to 1/2
9/12 = (3x3)/(3x4) = 3/4 isn’t equal to 1/2
7/14 = (7x1)/(7x2) = 1/2

Thus, 6/12, 15/30, 5/10 and 7/14 are equal to 1/2